# Discovering Properties of Inflectional Morphology in Neural Emergent Communication

### v1

@@ Appendix D HASLen for Natural and Emergent Languages
Appendix E Segmentation Analysis of Natural Languages
@@
Appendix F Agent Implementation Details and Task Accuracy
In line with prior related works, we implement the sender and receiver agents as single layer gated recurrent unit (GRU; Cho et al., 2014) models using the EGG framework Kharitonov et al. (2019), with hidden dimensions of 500. Observed states were implemented by concatenating one-hot vectors for each attribute. Senders are optimized using REINFORCE Williams (1992) and receivers with backpropagation following Chaabouni et al. (2019) using the accuracy (i.e., proportion of attributes correctly reconstructed) as a loss, with cross-entropy loss as a supplementary loss term. It was found in initial experiments that this allowed smoother and more reliable convergence with lower sequence lengths, which is desirable for obtaining interpretable and minimally complex languages. Each run occurred on a single A40 GPU with 48GB VRAM and took about 8 hours, although many models converged before then. Models generally converged around the 400 epoch mark.
We ran experiments with agents 8 times for each setting with different random initializations. In the default Attr-Val setting, 7 of 8 emerged languages were successful, i.e., achieved greater than 90% accuracy on the reconstruction task. For other experiments, including with the articulation pressure discussed in Section 6.1 and the inflectional Attr-Val setting, all languages were considered successful.
Appendix G Ease-of-Learning Pressure Experiment
@@
Appendix H Articulation Pressure Details
@@
Appendix I Additional Notes On Inflection Setting
| Setting | Articulation? | BPELen_96 | F-TopSim | HAS BoSDis | BPE BoSDis |
| 16x16x16 | Y | 3.534 | -0.070 | 186 | 12.1 |
| N | 3.825 | 0.0112 | 3.25 | 1.25 | |
| 42x2x3 | Y | 2.340 | -0.0083 | 1.56 | 30.7 |
| N | 3.257 | -0.0078 | 113 | 45.4 |
Appendix J Concatenativity and Fusionality Metrics for all EmCom Experiments
In Table 7, we summarize the concatenativity and ease of fusion metrics for each combination of Attr-Val setting and presence of the articulation pressure. As observed there, the average F-TopSim values are universally close to 0 in emergent languages, indicating some inconsistent degree of fusion. Only in the default setting without articulation pressure is F-TopSim positive on average, however, indicating a high degree of fusion. A t-test for mean F-TopSim between articulation pressure and no articulation pressure in the default setting yields highly significant results (p << 0.05). A possible explanation for this is that fusion creates more total symbols, and articulation pressure discourages extra symbols, as there are only so many readily distinguishable symbols that abide by the articulation rule.

### v2

@@ Appendix F HASLen for Natural and Emergent Languages
Appendix G Segmentation Analysis of Natural Languages
@@
Appendix H Ease-of-Learning Pressure Experiment
@@
Appendix I Articulation Pressure Details
@@
Appendix J Additional Natural Language Comparisons
Figure 10 shows BPE compression computed for Spanish, Arabic, emergent languages without articulation pressure (Em. Base) and emergent languages with articulation pressure (Em. Artic.). Figure 11 shows TopSim ranges for the same experiments. Due to lack of a sufficiently large grammatical feature in our Arabic dataset, we were unable to compute an Arabic baseline for the 42 6 condition.
Appendix K License Information
We provide license information and links for the following artifacts used in this work:
EGG Framework: MIT License.
Fred Jehle’s Spanish Verb Dataset: Creative Commons Attribution-NonCommercial-ShareAlike 3.0 Unported License.
Arabic-Urdu Conjugation Dataset